OpenStudy (anonymous):

Write the equation in its equivalent logarithmic form 62 = 36

Parth (parthkohli):

\( \color{Black}{\Rightarrow \log_ab = x \Longrightarrow a^x = b }\)

OpenStudy (mertsj):

But 62 does NOT equal 36

Parth (parthkohli):

@Mertsj lol he meant 6^2 = 36

OpenStudy (anonymous):

Yeah . ^

OpenStudy (mertsj):

@ParthKohli I know but perhaps he could learn to post correctly.

OpenStudy (anonymous):

Perhaps you could just help me out .

OpenStudy (mertsj):

Instead of expecting the rest of us to infer meaning to something meaningless.

Parth (parthkohli):

He copied it from some site because when you copy \(6^2\), you'll get 62.

OpenStudy (mertsj):

The log is the exponent. So we begin with log = 2

Parth (parthkohli):

What are a,b and x here?

OpenStudy (mertsj):

So what is the base?

OpenStudy (mertsj):

Well, it is clearly 6

OpenStudy (mertsj):

So now we can say: \[\log_{6}( )=2 \]

OpenStudy (mertsj):

Clearly 6^2=36 and so we write: \[\log_{6}36=2 \]

Parth (parthkohli):

That's it